OpenStudy (anonymous):

A rectangle has an area of 72 square units. The width of the rectangle is 9 units. The length of the rectangle is 2x + 4. What is the rectangle's length?

OpenStudy (anonymous):

the area of a rectangle is length times width. A = L * W 72 = (2x+4) * 9 8 = 2x + 4 4 = 2x x = 2 L = 2x + 4 = 4 + 4 = 8